OpenStudy (anonymous):

find the inverse of the function defined by f(x) = 5^x-2 +3 ..

OpenStudy (anonymous):

find the inverse of the function defined by \[f (x) = 5^{x-2} +3\]

OpenStudy (anonymous):

whats the first thing u think u need to do before i get started

OpenStudy (anonymous):

i think i do first the steps in inverse function \[ f(x)=5^{x-2}+3\] then \[y=5^{x-2}+3\] then \[y-3=5^{x-2}\] then \[y-3 = \log (x-2)\] i dont know whats next.. help me :D

OpenStudy (anonymous):

very good

OpenStudy (anonymous):

multiply both sides by x-2

OpenStudy (anonymous):

then distrubute

OpenStudy (anonymous):

how??

hartnn (hartnn):

you're correct till \(y-3 = 5^{x-2}\) now take log on BOTH sides.

OpenStudy (anonymous):

when it comes to log, i dont know.. help!

Parth (parthkohli):

Can you simplify \(\log(5^{x - 2})\)?

hartnn (hartnn):

ok, \( \log (y-3) = \log (5^{x-2})\) for the right side, use the property that \(\log A^B = B \log A\)

hartnn (hartnn):

still not clear ?

OpenStudy (anonymous):

srry step out really fast

OpenStudy (anonymous):

wait.. still thinking :D

hartnn (hartnn):